### v6

@@ 4.1 Preliminary lemmas
    $\dot{H}^{\gamma,q}$ is continuously embedded in $(L^{q},\dot{H}^{1,q})_{\beta,q}$;
@@
    $(L^{q},\dot{H}^{1,q})_{\gamma,q}$ is continuously embedded in $\dot{H}^{\beta,q}$.
@@
Let $0\leqslant\alpha\leqslant\beta<\infty$ and $q\in(1,\infty)$. Then for all $t\in(0,T]$ we have
@@
Let $p,q\in(1,\infty)$. Then for any $g\in L^{p}(0,T;L^{q})$, the following maximal $L^{p}$-regularity estimate holds:
$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0}g}\Big{\rVert}_{L^{p}(0,T;L^{q})}+\lVert{S_{0}g}\rVert_{L^{p}(0,T;\dot{H}^{2,q})}+\lVert{S_{0}g}\rVert_{C([0,T];(L^{q},\dot{H}^{2,q})_{1-1/p,p})}$
[…]
@@
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, the following stochastic maximal $L^{p}$-regularity estimate holds:
$\lVert{S_{1}g}\rVert_{L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1}g}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}^{1,q})}$
[…]
For any $\alpha\in\mathbb{R}$ and $q\in(1,\infty)$, let $\dot{H}_{h}^{\alpha,q}$ denote the space $X_{h}$ equipped with the norm
$\lVert{v_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}}:=\lVert{(-\Delta_{h})^{\alpha/2}v_{h}}\rVert_{L^{q}},\quad\forall v_{h}\in X_{h}.$
@@
Let $p,q\in(1,\infty)$. Then for any $g_{h}\in L^{p}(0,T;X_{h})$, the following discrete maximal $L^{p}$-regularity estimate holds:
$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0,h}g_{h}}\Big{\rVert}_{L^{p}(0,T;\dot{H}_{h}^{0,q})}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\\ \leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
Lemma 4.6.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,X_{h}))$, the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
[…]
Lemma 4.7.
The following assertions are valid:
  1. (i)
    For all $q\in[1,\infty]$, it holds that
    $\lVert{P_{h}}\rVert_{\mathcal{L}(L^{q})}\text{ is uniformly bounded with respect to $h$}.$ (4.2)
  2. (ii)
    For all $\alpha\in[0,2]$ and $q\in(1,\infty)$, $P_{h}$ has the following approximation property:
    $\lVert{I-P_{h}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha}.$ (4.3)
  3. (iii)
    For all $0\leqslant\alpha<\beta\leqslant 2$ and $q\in(1,\infty)$, the following inequality holds:
    $\lVert{\Delta_{h}P_{h}-P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{\beta,q},\dot{H}_{h}^{\alpha-2,q})}\leqslant ch^{\beta-\alpha}.$ (4.4)
  4. (iv)
    For all $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t>0$, the following inequalities are satisfied:
    $\lVert{e^{t\Delta_{h}}}\rVert_{\mathcal{L}(\dot{H}_{h}^{0,q},\dot{H}_{h}^{\alpha,q})}\leqslant ct^{-\alpha/2},$ (4.5)
    $\lVert{e^{t\Delta_{h}}P_{h}-P_{h}e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}+\lVert{(I-P_{h})e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}$
    $\leqslant c\min\Big{\{}h^{\alpha},\frac{h^{2}}{t^{1-\alpha/2}}\Big{\}}.$ (4.6)
Remark 4.1.
Lemma 4.1 is established by Proposition 1.4 from [21] and Theorem C.4.1 in [10]. The inequality asserted in Lemma 4.2 is a direct consequence of Theorem 2.12 and Equation (2.128) as presented in [33]. The proof of Lemma 4.3 is provided in [32] and Corollary 1.14 from [21]. The proof of Lemma 4.4 is detailed in [30]. The proof of Lemma 4.5 is accessible in [15], and the proof of Lemma 4.6 can be found in [16]. The assertions in Lemma 4.7 are widely recognized and can be found in standard references such as [6, 28].
Lemma 4.8.
For any $q\in(1,\infty)$, the following inequalities hold:
$\lVert{F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),$ $\forall v\in L^{q};$ (4.7)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}),$ $\forall v\in\dot{H}^{1,q};$ (4.8)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{s_{1},q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{s_{2},q}}),$ $\forall v\in\dot{H}^{s_{2},q},\,\forall 0\leqslant s_{1}<s_{2}<1;$ (4.9)
$\lVert{F(u)-F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},$ $\forall u,v\in L^{q}.$ (4.10)
@@
We only prove Eq. 4.9, the proof of Eqs. 4.7, 4.8 and 4.10 being easier. Let $\Omega_{\gamma}$ denote a probability space on which a sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of independent standard Gaussian random variables is defined, and let $\mathbb{E}_{\gamma}$ denote the expectation operator with respect to $\Omega_{\gamma}$. The rest of the proof is divided into two steps.
@@
For any $v\in\dot{H}^{1,q}$, we infer that
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\Big{(}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}\mu\Big{)}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\Big{(}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q}$
$=c\Big{(}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\Big{(}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}(1+\lvert{v}\rvert+\lvert{\nabla v}\rvert)^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q},$
where step $(\mathrm{i})$ invokes Theorem 6.2.6 from [11], and step $(\mathrm{ii})$ relies on Lipschitz conditions Eqs. 2.3 and 2.4. Using the summability condition Eq. 2.5 and the standard inequality
@@
Analogously, employing Eq. 2.2 yields
@@
and utilizing Eq. 2.3 results in
@@
This result, when combined with the continuous embedding property
@@
yields
@@
Applying Theorem 6.2.6 from [11] once again, we arrive at the assertion in Eq. 4.11.
@@
$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}=\lVert{\lim_{N\to\infty}A_{N}v}\rVert_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})},$
which stems directly from the definition of $F$ as stated in (2.6) and is justified by Theorem 9.1.19 in [11]. By invoking the inequality (4.11), we apply Fatou’s Lemma to deduce that
@@
As established by Theorem 4.36 in [21], the interpolation space $(L^{2},\dot{H}^{1,q})_{s,2}$ coincides with $\dot{H}^{s,2}$, sharing equivalent norms across the interval $s\in(0,1)$. Consequently, we refine the estimate Eq. 4.9 to the following form:
$\|F(v)\|_{\gamma(H,\dot{H}^{s,2})}\leqslant c(1+\|v\|_{\dot{H}^{s,2}}),\quad\forall v\in\dot{H}^{s,2},\,\forall s\in(0,1).$
@@
It is a well-established fact that the space $L^{2}$ admits an orthonormal basis $\{\phi_{n}\}_{n\in\mathbb{N}}$, comprised of eigenvectors of the negative Laplace operator $-\Delta$. Furthermore, this basis $\{\phi_{n}\}_{n\in\mathbb{N}}$ forms a complete orthogonal system within the space $\dot{H}^{s,2}$ for all $s\in(0,\infty)$.
@@
$\mathrm{d}y_{n}(t)=\big{(}\Delta y_{n}+y_{n}-y_{n}^{3}\big{)}(t)\,\mathrm{d}t+F(y_{n}(t))\,\mathrm{d}W_{H}(t),\quad 0<t\leqslant T,$ (4.12a)
@@
Step 1. Let $G_{n}(t):=e^{t\Delta}v_{n}$, $t\in[0,T]$. Since $v_{n}\in\dot{H}^{2,q}$ for each $q\in(2,\infty)$, by Lemma 4.2 with $\alpha=\beta=2$, we have
@@
In view of Eq. 4.8 and the fact that $v_{n}\in\dot{H}^{1,2}$, we use [17, Theorem 1.1] to deduce that Eq. 4.12 has a unique mild solution $y_{n}$ and
@@
which implies $y_{n}-y_{n}^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for each $p\in(2,\infty)$. Hence, using Lemma 4.3 gives
@@
Since Eqs. 4.8 and 4.15 imply that $F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for each $p\in(2,\infty)$, by Lemma 4.4 we obtain
@@
$y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n}),$ (4.18)
and combining Eqs. 4.14, 4.16 and 4.17, we conclude that
@@
From the continuous embeddings of $\dot{H}^{2,2}$ into $\dot{H}^{1,6}$ and $\dot{H}^{2,2}$ into $L^{q}$ for all $q\in(2,\infty)$, it follows that
@@
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$ (4.20)
Combining Eqs. 4.14, 4.18, 4.19 and 4.20 yields
@@
thereby implying, through the continuous embedding of $\dot{H}^{2,6}$ into $\dot{H}^{1,q}$ for all $q\in(2,\infty)$, that
@@
This regularity result allows us to enhance Eq. 4.17 to
@@
which, combined with Eqs. 4.14, 4.18 and 4.19, implies
@@
$\mathbb{E}\Big{(}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}-\lVert{v_{n}}\rVert_{L^{q}}^{p}\Big{)}$
@@
$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lambda_{k}\lvert{f_{k}(y_{n}(s))}\rvert^{2}\,\mathrm{d}\mu\,\mathrm{d}s$
$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)f_{k}(y_{n}(s))\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s$
@@
$I_{1}\leqslant p\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s.$
Regarding $I_{2}+I_{3}$, by using Hölder’s inequality in conjunction with Young’s inequality, and employing the summability condition Eq. 2.5 alongside the growth condition Eq. 2.2, a straightforward calculation leads to the following bound for $I_{2}+I_{3}$:
$I_{2}+I_{3}\leqslant c\,\mathbb{E}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s.$
@@
$\mathbb{E}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{v_{n}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s,\quad\forall t\in[0,T].$
@@
$\lVert{y_{n}(t)}\rVert_{L^{p}(\Omega;L^{q})}\leqslant c\big{(}1+\lVert{v_{n}}\rVert_{L^{q}}\big{)},\quad\forall t\in[0,T],$
@@
Step 1. Let us prove that Eq. 1.1 admits a unique mild solution $y$ and Eq. 3.1 holds. Let $p,q\in(2,\infty)$ be arbitrary but fixed. For each $n\in\mathbb{N}$, let $y_{n}$ be the mild solution of Eq. 4.12 with $v_{n}$ defined by
@@
For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y_{m}-y_{n}$. By definition, we infer that, for any $t\in[0,T]$,
$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+y_{n}^{3}-y_{m}^{3})(t)\,\mathrm{d}t+\big{(}F(y_{m}(t))-F(y_{n}(t))\big{)}\,\mathrm{d}W_{H}(t).$
@@
$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{v_{m}-v_{n}}\rVert_{L^{3q}}.$
On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $\{\phi_{n}\}_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $\{v_{n}\}_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$, converging to $v$. Consequently, there exists a unique $y\in L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})$ such that
@@
For any $n\in\mathbb{N}$, application of Hölder’s inequality yields
@@
Using Lemma 4.3, this directly implies the following convergence for $S_{0}(y_{n}-y_{n}^{3})$:
@@
Furthermore, invoking Eqs. 4.10 and 4.22 yields
@@
where $G$ is defined by Eq. 2.11, and $G_{n}(t):=e^{t\Delta}v_{n}$ for each $t\in[0,T]$. Finally, combining the derived convergences of $y_{n}$, $G_{n}$, $S_{0}(y_{n}-y_{n}^{3})$, and $S_{1}F(y_{n})$, and recalling that $y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n})$, by passing to the limit $n\to\infty$ we obtain

### v7

@@ 4.1 Preliminary Lemmas
It is a well-established result that for any $q\in(1,\infty)$ and $s\in(0,2)$, the space $\dot{H}^{s,q}$ coincides with the complex interpolation space $[L^{q},\dot{H}^{2,q}]_{s/2}$, equipped with equivalent norms. This result is documented, for example, in Chapter 16 of [38]. The combination of Proposition 1.4 from [22] and Theorem C.4.1 in [12] yields the following continuous embeddings.
@@
    The space $\dot{H}^{\gamma,q}$ is continuously embedded in $(L^{q},\dot{H}^{1,q})_{\beta,q}$;
@@
    The space $(L^{q},\dot{H}^{1,q})_{\gamma,q}$ is continuously embedded in $\dot{H}^{\beta,q}$.
Theorem 2.12 and inequality (2.128) in [38] provide the following standard estimate for the growth of the analytic semigroup generated by the Laplace operator with homogeneous Dirichlet boundary conditions.
@@
Let $0\leqslant\alpha\leqslant\beta<\infty$ and $q\in(1,\infty)$. For all $t\in(0,T]$, the following estimate holds:
@@
For the deterministic convolution operator $S_{0}$, defined by Eq. 2.7, we have the following celebrated deterministic maximal $L^{p}$-regularity estimate; see Theorem 6.1.11 in [30] and Corollary 1.14 from [22].
@@
Let $p,q\in(1,\infty)$. Then for any $g\in L^{p}(0,T;L^{q})$, the following inequality holds:
$\big{\lVert}\frac{\mathrm{d}}{\mathrm{d}t}S_{0}g\big{\rVert}_{L^{p}(0,T;L^{q})}+\lVert{S_{0}g}\rVert_{L^{p}(0,T;\dot{H}^{2,q})}+\lVert{S_{0}g}\rVert_{C([0,T];(L^{q},\dot{H}^{2,q})_{1-1/p,p})}\leqslant c\lVert{g}\rVert_{L^{p}(0,T;L^{q})}.$
For the stochastic convolution operator $S_{1}$, defined by Eq. 2.8, we present the following fundamental stochastic maximal $L^{p}$-regularity estimate; see Theorems 1.1 and 1.2 in [26].
@@
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, the following inequality holds:
$\lVert{S_{1}g}\rVert_{L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1}g}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}^{1,q})}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}.$
The operator $F$, defined by Eq. 2.6, possesses the following growth and Lipschitz continuity properties.
@@
Let $q\in(1,\infty)$. The following inequalities are valid:
$\lVert{F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\left(1+\lVert{v}\rVert_{L^{q}}\right),$ for all $\text{for all }v\in L^{q};$ (4.1)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c\left(1+\lVert{v}\rVert_{\dot{H}^{1,q}}\right),$ for all $\text{for all }v\in\dot{H}^{1,q};$ (4.2)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{s_{1},q})}\leqslant c\left(1+\lVert{v}\rVert_{\dot{H}^{s_{2},q}}\right),$ for $\text{for all }v\in\dot{H}^{s_{2},q},\text{ where }0\leqslant s_{1}<s_{2}<1;$ (4.3)
$\lVert{F(u)-F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},$ for all $\text{for all }u,v\in L^{q}.$ (4.4)
@@
We focus our demonstration on Eq. 4.3; the proofs of Eq. 4.1, Eq. 4.2, and Eq. 4.4 follow similar, yet more straightforward, lines of reasoning. For the definition of $\gamma$-radonifying operators, we direct the reader to Section 9.1 of [13]. Consider $\Omega_{\gamma}$ as a probability space supporting a sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of independent standard Gaussian random variables. Let $\mathbb{E}_{\gamma}$ denote the expectation operator with respect to the probability space $\Omega_{\gamma}$. The subsequent argument is structured in two main steps.
@@
For any $v\in\dot{H}^{1,q}$, we deduce that
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\bigg{[}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}\mu\bigg{]}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q}$
$=c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}(1+\lvert{v}\rvert+\lvert{\nabla v}\rvert)^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q},$
where step $(\mathrm{i})$ uses the famous Kahane-Khintchine inequality (see Theorem 6.2.6 from [13]), and step $(\mathrm{ii})$ relies on Lipschitz conditions Eqs. 2.3 and 2.4. Using the summability condition Eq. 2.5 and the standard inequality
@@
Analogously, employing the growth condition Eq. 2.2 yields
@@
and utilizing the Lipschitz continuous condition Eq. 2.3 results in
@@
Combining this result with the continuous embedding property
@@
which can be easily verified by definition, yields
@@
Applying the Kahane-Khintchine inequality once again, we arrive at the assertion in Eq. 4.5.
@@
$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}=\Big{\lVert}\lim_{N\to\infty}A_{N}v\Big{\rVert}_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})},$
which follows directly from the definition of $F$ as stated in (2.6) and is justified by Theorem 9.1.19 in [13]. Using the inequality (4.5), we apply Fatou’s Lemma to deduce that
@@
As established by Theorem 4.36 in [22], the interpolation space $(L^{2},\dot{H}^{1,2})_{s,2}$ coincides with $\dot{H}^{s,2}$, sharing equivalent norms for all $s\in(0,1)$. Consequently, we refine the inequality Eq. 4.3 to the following form:
$\|F(v)\|_{\gamma(H,\dot{H}^{s,2})}\leqslant c(1+\|v\|_{\dot{H}^{s,2}}),\quad\text{for all $v\in\dot{H}^{s,2}$, where $s\in(0,1)$}.$
@@
It is well known that the space $L^{2}$ admits an orthonormal basis $(\phi_{n})_{n\in\mathbb{N}}$, consisting of eigenvectors of the negative Laplace operator $-\Delta$ with homogeneous Dirichlet boundary conditions. This basis $(\phi_{n})_{n\in\mathbb{N}}$ forms a complete orthogonal system in the space $\dot{H}^{s,2}$ for all $s\in(0,\infty)$.
@@
$\mathrm{d}y_{n}(t)=\big{(}\Delta y_{n}+y_{n}-y_{n}^{3}\big{)}(t)\,\mathrm{d}t+F(y_{n}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (4.6a)
@@
Step 1. Define $G_{n}(t):=e^{t\Delta}v_{n}$ for $t\in[0,T]$. Given that $v_{n}\in\dot{H}^{2,q}$ for every $q\in(2,\infty)$, it follows from Lemma 4.2 with $\alpha=\beta=2$ that
@@
In view of Eq. 4.2 and the fact that $v_{n}\in\dot{H}^{1,2}$, we use [18, Theorem 1.1] (see also, [21, Lemma 2.3 and Theorem 3.1]) to deduce that Eq. 4.6 has a unique mild solution $y_{n}$ and
@@
which in turn implies that $y_{n}-y_{n}^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for all $p\in(2,\infty)$. Applying Lemma 4.3, we obtain
@@
Since (4.2) and (4.9) ensure that $F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for every $p\in(2,\infty)$, Lemma 4.4 allows us to conclude that
@@
$y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n}),$
and combining (4.8), (4.10), and (4.11), we deduce that
@@
The continuous embeddings of $\dot{H}^{2,2}$ into $\dot{H}^{1,6}$ and $\dot{H}^{2,2}$ into $L^{q}$ for all $q\in(2,\infty)$ yield
@@
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$
Combining these refined regularity results with (4.8), we arrive at
@@
which, through the continuous embedding of $\dot{H}^{2,6}$ into $\dot{H}^{1,q}$ for all $q\in(2,\infty)$, implies that
@@
This regularity result allows us to further improve (4.11) to
@@
which, in conjunction with (4.8) and (4.12), yields
@@
$\mathbb{E}\Big{[}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}-\lVert{v_{n}}\rVert_{L^{q}}^{p}\Big{]}$
@@
$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\Biggl{[}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lambda_{k}\lvert{f_{k}(y_{n}(s))}\rvert^{2}\,\mathrm{d}\mu\,\mathrm{d}s\Biggr{]}$
$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\Biggl{\{}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\lambda_{k}\bigg{[}\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)f_{k}(y_{n}(s))\,\mathrm{d}\mu\bigg{]}^{2}\,\mathrm{d}s\Biggr{\}}$
@@
$I_{1}\leqslant p\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
For $I_{2}+I_{3}$, using Hölder’s inequality in conjunction with Young’s inequality, and employing the summability condition Eq. 2.5 along with the growth condition Eq. 2.2, we derive the following bound:
$I_{2}+I_{3}\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
@@
$\mathbb{E}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{v_{n}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]},\quad\forall t\in[0,T].$
@@
$\lVert{y_{n}}\rVert_{C([0,T];L^{p}(\Omega;L^{q}))}\leqslant c\big{(}1+\lVert{v_{n}}\rVert_{L^{q}}\big{)},$
@@
Step 1. We proceed to show that equation Eq. 1.1 has a unique mild solution $y$, and that Eq. 3.1 is satisfied. Let $p,q\in(2,\infty)$ be fixed but arbitrary. For each $n\in\mathbb{N}$, denote by $y_{n}$ the mild solution of equation Eq. 4.6, where $v_{n}$ is defined as
@@
For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y_{m}-y_{n}$. By definition, we infer that
$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+y_{n}^{3}-y_{m}^{3})(t)\,\mathrm{d}t+\big{(}F(y_{m}(t))-F(y_{n}(t))\big{)}\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
@@
$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{v_{m}-v_{n}}\rVert_{L^{3q}}.$ (4.14)
On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $(\phi_{n})_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $(v_{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$, converging to $v$. Consequently, there exists a unique process $y\in L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})$ such that
@@
For any $n\in\mathbb{N}$, applying Hölder’s inequality yields
@@
Using Lemma 4.3, this directly implies the following convergence of $S_{0}(y_{n}-y_{n}^{3})$:
@@
Furthermore, invoking Eqs. 4.4 and 4.15 yields
@@
where $G$ is defined by Eq. 2.10, and $G_{n}(t):=e^{t\Delta}v_{n}$ for each $t\in[0,T]$. Finally, combining the derived convergences of $y_{n}$, $G_{n}$, $S_{0}(y_{n}-y_{n}^{3})$, and $S_{1}F(y_{n})$, and recalling that $y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n})$, we obtain by passing to the limit $n\to\infty$ that